Alliance Company’s budgets production of 25,000 units in January and 29,000 units in the February. Each finished unit requires 4 pounds of raw material K that costs $3.50 per pound. Each month’s ending raw materials inventory should equal 40% of the following month’s budgeted materials. The January 1 inventory for this material is 40,000 pounds. What is the budgeted materials cost for January?
Answer #1

Calculation of budgeted materials cost for January.

Raw material required for January budgeted production (A) = 25,000*4 pounds =100,000 pounds

Beginning inventory of raw material in January (B)= 40,000 pounds

Ending inventory of material for January (C) = 40% of the February month’s budgeted materials=29,000*4*40%=46,400

So  budgeted materials cost for January =[( A + C ) - B ] *$3.50

=100,000+46,400-40,000 *$3.50

=$372,400